# Learning to Deliberate: Meta-policy Collaboration for Agentic LLMs with Multi-agent Reinforcement Learning

### v1

@@ Introduction
As shown in Table 1, our deliberation-based multi-agent framework achieves the highest average accuracy (55.37%) across six diverse reasoning benchmarks, outperforming both debate-style methods (e.g., DyLAN 51.50%) and dynamic approaches (e.g., AgentPrune 51.60%). Unlike single-agent methods (e.g., Vanilla, CoT, SC), which lack inter-agent coordination, and debate-style models, which suffer from unstable decision-making and weak convergence, our framework jointly optimizes decision-making and consensus formation to enhance multi-agent reasoning. While routing methods attempt to improve efficiency through static interaction graphs and heuristic specialization, they remain limited by their inability to support dynamic deliberation or mutual influence during execution. In contrast, our approach establishes a unified, meta-cognitive deliberation process that enables agents to collaborate iteratively toward shared goals, improving accuracy especially in high-complexity tasks such as AMC and AIME.
@@
To verify that our framework generalizes beyond a single model family, we conduct additional experiments using LLaMA3 and Qwen2.5 as representative backbones. Table 2 summarizes the results, demonstrating consistent and robust accuracy gains across all settings. On LLaMA-3B, the weakest backbone in our evaluation, accuracy improves from 68.84% (GPTSwarm) to 80.52%, indicating that the learned meta-policy effectively compensates for limited model capacity. Scaling up to LMA-8B still yields a positive margin, lifting the best baseline from 84.53% to 86.35%. Switching to Qwen models is particularly informative, as these checkpoints already benefit from extensive instruction tuning; Qwen-8B alone reaches 91.05%, yet deliberative training boosts it further to 94.47%; and Qwen-3B accuracy increases from 84.23% to 89.77%.
@@
Our framework transitions from supervised fine-tuning (SFT) to progressively more effective reinforcement learning (RL) strategies that enhance agent coordination. As shown in Table 1, even the SFT-only variant performs competitively (52.09%), highlighting the strength of deliberation-based reasoning. However, it lacks outcome-driven feedback, limiting its ability to refine decision-making. Vanilla PPO performs worse (50.85%) due to unstable credit assignment and poor convergence in our multi-agent setting. As shown in Figure 3, PPO struggles especially on harder tasks (e.g., -3.62% on AMC), where its value-based updates fail to provide stable guidance. GRPO mitigates this by using group-relative ranking, stabilizing updates through peer comparison and improving performance on tasks like GSM8K (+1.97%) and MATH (+2.18%), but its reliance on raw reward magnitudes leads to sensitivity under noisy or heavy-tailed feedback. With reinforcement signals aligned to the group’s preference order, SoftRankPO yields the best overall results (55.37%) and more stable convergence dynamics (Figure 3).
@@
We conduct an ablation where the reward signal is rescaled by exponential factors $\{0.1,1.0,10.0\}$. As shown in Figure 5, GRPO suffers from noticeable instability: different scales induce diverging learning curves, and all variants exhibit a pronounced drop in performance after early gains. This behavior stems from GRPO’s reliance on perstate $z$-score advantages, where an outlier can dominate the local mean and variance, leading to unstable gradient updates. In contrast, SoftRankPO is robust across reward scales. Its accuracy consistently improves and stabilizes around the same level regardless of the reward scaling, with visibly smoother and more stable curves. This empirical observation aligns with our analysis in Theorem 3. By inducing a zero-mean, bounded-variance advantage shaped by standard Gaussian, SoftRankPO effectively decouples learning dynamics from the scale of rewards.
@@
As shown in Figure 7, SoftRankPO induces a significant behavioral shift: the PERSIST rate increases from 19.1% to 78.8%, while REFINE and CONCEDE jointly fall from 81% to 21%. This stark transition suggests that agents have learned to recognize and trust the high-quality outputs produced by the pretrained LLM. It reflects an emerging ability for self-assessment and confidence calibration. The resulting policy aligns with the principle of minimal intervention: instead of revising reflexively or adopting peer responses blindly, agents now act selectively, intervening only when they perceive a meaningful improvement. Such a pattern is directly encouraged by our reward shaping, which emphasizes marginal utility and penalizes redundant edits. Beyond accuracy, this leads to reduced token use (Table 3). Furthermore, a detailed case study is presented in Appendix C.

### v2

@@ Introduction
As shown in Table 1, our deliberation-based multi-agent framework achieves the highest average accuracy (55.74%) across six diverse reasoning benchmarks, outperforming both debate-style methods (e.g., DyLAN 51.50%) and dynamic approaches (e.g., AgentPrune 51.60%). Unlike single-agent methods (e.g., Vanilla, CoT, SC), which lack inter-agent coordination, and debate-style models, which suffer from unstable decision-making and weak convergence, our framework jointly optimizes decision-making and consensus formation to enhance multi-agent reasoning. While routing methods attempt to improve efficiency through static interaction graphs and heuristic specialization, they remain limited by their inability to support dynamic deliberation or mutual influence during execution. In contrast, our approach establishes a unified, meta-cognitive deliberation process that enables agents to collaborate iteratively toward shared goals, improving accuracy especially in high-complexity tasks such as AMC and AIME.
@@
To verify that our framework generalizes beyond a single model family, we conduct additional experiments using LLaMA3 and Qwen2.5 as representative backbones. Table 2 summarizes the results, demonstrating consistent and robust accuracy gains across all settings. On LMA-3B, the weakest backbone in our evaluation, accuracy improves from 68.84% (GPTSwarm) to 80.52%, indicating that the learned meta-policy effectively compensates for limited model capacity. Scaling up to LMA-8B still yields a positive margin, lifting the best baseline from 84.53% to 86.35%. Switching to Qwen models is particularly informative, as these checkpoints already benefit from extensive instruction tuning; Qwen-7B alone reaches 91.05%, yet deliberative training boosts it further to 94.47%; and Qwen-3B accuracy increases from 84.23% to 89.77%.
@@
Our framework transitions from supervised fine-tuning (SFT) to progressively more effective reinforcement learning (RL) strategies that enhance agent coordination. As shown in Table 1, even the SFT-only variant performs competitively (52.09%), highlighting the strength of deliberation-based reasoning. However, it lacks outcome-driven feedback, limiting its ability to refine decision-making. Vanilla PPO performs worse (50.85%) due to unstable credit assignment and poor convergence in our multi-agent setting. As shown in Figure 3, PPO struggles especially on harder tasks (e.g., -3.62% on AMC), where its value-based updates fail to provide stable guidance. GRPO mitigates this by using group-relative ranking, stabilizing updates through peer comparison and improving performance on tasks like GSM8K (+1.97%) and MATH (+2.18%), but its reliance on raw reward magnitudes leads to sensitivity under noisy or heavy-tailed feedback.
@@
We conduct an ablation where the reward signal is rescaled by exponential factors $\{0.1,1.0,10.0\}$. As shown in Figure 5, GRPO suffers from noticeable instability: different scales induce diverging learning curves, and all variants exhibit a pronounced drop in performance after early gains. This behavior stems from GRPO’s reliance on perstate $z$-score advantages, where an outlier can dominate the local mean and variance, leading to unstable gradient updates. In contrast, SoftRankPO is robust across reward scales. Its accuracy consistently improves and stabilizes around the same level regardless of the reward scaling, with visibly smoother and more stable curves. This empirical observation aligns with our theoretical analysis. By inducing a zero-mean, bounded-variance advantage shaped by standard Gaussian, SoftRankPO effectively decouples learning dynamics from the scale of rewards.
@@
As shown in Figure 7, SoftRankPO induces a significant behavioral shift: the PERSIST rate increases from 19.1% to 78.8%, while REFINE and CONCEDE jointly fall from 81% to 21%. This stark transition suggests that agents have learned to recognize and trust the high-quality outputs produced by the pretrained LLM. It reflects an emerging ability for self-assessment and confidence calibration. The resulting policy aligns with the principle of minimal intervention: instead of revising reflexively or adopting peer responses blindly, agents now act selectively, intervening only when they perceive a meaningful improvement. Such a pattern is directly encouraged by our reward shaping, which emphasizes marginal utility and penalizes redundant edits. Beyond accuracy, this leads to reduced token use (Table 3).
@@ Conclusion
Acknowledgments
This work was supported in part by a grant from the USC–Capital One Center for Responsible AI and Decision Making in Finance (CREDIF).